.

TimestampEvent TimeAnnotationPersonEvent TimePerson

.

15/02/2013FILES FOR THE CELLS #2 ARE LOCATED HERE: https://docs.google.com/folder/d/0B0yO8n6-0MjNU2tXSGxaT0NscXc/editMathieu Valat2013-02-15

.

05/12/2012 03:24:032012-11-29-17:00CETCell uses wire of constantan with oxide coating: 0.2mm diam, 1m long, 16Ω.
Loaded with 2.62b of H2 then topped with Ar up to 3.5b.
15 steps are 1 hour long (28.5W, 38, 42.5, 47.4, 52.3, 57.5, 62, 72.66, 76,23, 82.6, 89.23, 94.79, 100.52, 109.16W)
File is Cell#2-Calib-20121129-75H2-25Ar-P3.5.xlsx
Mathieu Valat2013-02-15

.

05/12/2012 05:29:042012-11-30-09:30CETUnloaded from 3.4b to 2b of H2/Ar.
File is Cell#2-Calib-20121130-75H2-25Ar-P2.xlsx
Mathieu Valat2012-12-05

.

05/12/2012 05:32:242012-12-01-00:30CETUnloaded from 1.94b to 1b of H2/Ar.
File is Cell#2-Calib-20121201-75H2-25Ar-P1.xlsx
Mathieu Valat2012-12-05

.

05/12/2012 05:43:182012-12-01-17:00CETObjective was to roughly define Power vs T_mica for constant pressure (1b) under H2/Ar. After the power is set I wait until 80% of temperature rise then I compensate the pressure gain. When apparent stabilization for T Mica is reached I re compensate the gain of pressure. Then I change the power and redo this operations. P_in were 40W, 50W, 60W, 70W, 80W.
File is Cell#2-Calib-20121201-75H2-25Ar-P1constant.xlsx
Mathieu Valat2012-12-05

.

05/12/2012 08:16:142012-12-01-20:00CETLoaded with H2@1b.
File is Cell#2-Calib-20121202-H2-P1.xlsx
Mathieu Valat2012-12-05

.

05/12/2012 08:21:272012-12-02-10:30CETLoaded with H2@0.5b.
File is Cell#2-Calib-20121202-H2-P0.5.xlsx
Mathieu Valat2012-12-05

.

05/12/2012 08:26:202012-12-03-11:30CETVacuuming down the chamber and heating the wire to 1.9A during 3 hours.
Objective is to release H2 from wire.
Had to reduce to 1.5A quickly because it was blowing red due to lack of thermal convection.
File is Cell#2-Calib-20121203-He-baking1.5A.xlsx
Mathieu Valat2012-12-10

.

05/12/2012 08:32:362012-12-03-19:40CETLoaded with He@3.5b.
File is Cell#2-Calib-20121203-He-P3.5.xlsx
Mathieu Valat2012-12-05

.

05/12/2012 08:35:572012-12-04-10:30CETLoaded with He@2b.
File is Cell#2-Calib-20121204-He-P2.xlsx
Mathieu Valat

.

05/12/2012 08:38:432012-12-05-09:20CETLoaded with He@1b.
hugnet_router failed and stopped reporting data before and during lunch.
Mathieu Valat

.

05/12/2012 08:55:332012-12-05-15:10CETHe@1b from previous loading.
File is Cell#2-Calib-20121205-He-P1.xlsx
Mathieu Valat

.

2012-12-06-09:00CETHe@0.5b from previous loading.
File is Cell#2-Calib-20121206-He-P0.5.xlsx
Mathieu Valat2012-12-06

.

2012-12-07-09:00CETFlushed with 1.5b of H2 then done gaz mixture (75%H2+25%Ar) File is Cell#2-Calib-20121207-75H2-25Ar-P3.5.xslxMathieu Valat2012-12-08

.

2012-12-08-11:30CETMimic run of the future operating parameters (constant 48W w/75%H2+25%Ar@2b). Power slowly set to 48W to see how the wire resistivity vary at low power, this run should last 30h. Will help define leakage speed. File is Cell#2-Calib-20121208-75H2-25Ar-P2-48W.xslxMathieu Valat2012-12-08

.

2012-12-09-20:00CETBaked the cell under vacuum to release H for the internal part. Temperature of the wire went too high and it started vaporizing inside the cell. Had to change the tube with a spare one (same original tube). Cell is open and the wire is changed with the active wire. Flushed 4 time with He 3.5b after mounting. Calibration is done up to 1.9A to avoid changing the material structure before the run. Increased slowly the power manually to plot the transitionnal change in imperance at begining. Then calibration script run. File is Cell#2-360L-20121209-He-P3.5.xlsxMathieu Valat2012-12-10

.

2012-12-10-09:00CETPressure set to 2 bar. File is Cell#2-360L-20121210-He-P2.xlsxMathieu Valat2012-12-10

.

2012-12-10-16:00CETPressure set to 1 bar. File is Cell#2-360L-20121210-He-P1.xlsxMathieu Valat2012-12-10

.

2012-12-11-09:00CETPressure set to 1 bar. File is Cell#2-360L-20121211-He-P0.5.xlsxMathieu Valat2012-12-11

.

2012-12-11-15:00CETVacuum down the chamber. Wire the cell to the other power supplies. Use the first to put minimal amount of energy in the active wire to measure resistivity. Plug the second on to the NiCr wire. Flush and load with H2@2b. Impedance is 19.4 ohms. Increase the power to maximum voltage of 56V. Tmica rose to 165°C. R/Ro lost 10%, then I warp aliminium foid around the ends of better insulate and distribute the heat. But it acted like a disipator. I ended up let a sheet of aluminum foid fall on the cell to increase the temperature artificially to 175°C and reach maximum loading temperature extracted form Ryan's data. Picture: https://docs.google.com/open?id=0B9qCtGOFmvhmMXpOb2hsbTZKd1kMathieu Valat2012-12-12

.

2012-12-11-23:03CETI decided to let the cell run by itself and stop struggling with aluminum foid.Mathieu Valat2012-12-12

.

2012-12-12-09:00CETPower down. Let it cool. Impedance is 16.3 ohmsMathieu Valat2012-12-12

.

2012-12-12-10:00CETPower to the NiCr wire again to see if it absorb again. Strange variation of Celani wire at start. Reduced to minimal current in active wire few minutes before.Mathieu Valat2012-12-12

.

2012-12-12-11:20CETWiring the 2 power supplies in series on the nichrome wire to get more power. Reducing the pressure at 1 bar.Mathieu Valat2012-12-12

.

2012-12-12-14:08CETSlight decrease of Pin to confirm triggering affect of the temperatureMathieu Valat2012-12-12

.

2012-12-12-16:13CETLet the cell cool down, renew the H2.Mathieu Valat2012-12-12

.

12/12/2012 10:15:20Modified the data fields to include P_xs calculations. There is a P_out_Low based on the most conservative calibration line and a P_out_High based on the line from the first calibration which is what the unloaded Celani wire was close to under helium.Ryan Hunt

.

2012-12-12-19:00CETPowering the NiCr wire to check if the loading is optimal again. Then power the Celani wire to the 48W and see if there is a change in excess energy between indirect and direct heating.Mathieu Valat2012-12-12

.

2012-12-12-19:25CETResistivity was dropping at Tmica=130°C. Reducing Pin to get back there. Just in case...Mathieu Valat2012-12-12

.

2012-12-13-10:30CETLoading seems to reached maximum for this run (imperdance reached a plateau). Truning off the power and let it cool for one hour. Then I will restart again.Mathieu Valat2012-12-13

.

2012-12-13-11:40CETCell vacuumed then flushed with 3.5b of H2. Reloaded with 1b of pure H2. Direct power to 48WMathieu Valat2012-12-13

.

2012-12-14-11:30CETPower in set to 54W. I = 1.83AMathieu Valat2012-12-14

.

2012-12-15-12:00CETPower down. Let it cool. Setting gaz mixture 75%H2+25%Ar.Mathieu Valat2012-12-15

.

2012-12-16-17:20CETPower down. Cell is prepared with 1.5b of Ar for travel to Switzerland.Mathieu Valat2012-12-16

.

2012-12-17-17:20CETIntallation of Euro Cell in Switzerland labNicolas Chauvin2012-12-19

.

2012-12-18-15:00CETVacuum the cell to deload of Celani wire Nicolas Chauvin2012-12-19

.

2012-12-18-17:00CETAccident with NiChrome wire. Too much current broke the NiChrome wireMathieu Valat2012-12-19

.

2012-12-18-18:00CETPlan B: Continue deloading Celani wire by putting current in Celani wireNicolas Chauvin2012-12-19

.

2012-12-19-19:00CETCovered cell with copper sheet to improve thermalization inside the cell without driving too much current in the cell, still deloading the wire under vacuum.Nicolas Chauvin2012-12-19

.

2012-12-19-20:00CETInternet connection in the lab => back to live streamingNicolas Chauvin2012-12-19

.

2012-12-21-17:00CETStopped the vacuum pump. Put the cell at 1 atm of air. Stopped input power.Nicolas Chauvin2012-12-21

.

2013-01-10-19:00CETEuro cell reassermbled. Test vacuum for leaks.Nicolas Chauvin2013-01-10

.

2013-01-10-20:00CETCell loaded with 2.5 bar of Helium. Test for leaks. 20W in nichrome wireNicolas Chauvin2013-01-10

.

2013-01-11-11:00CETStopped power supply. Let the cell cool down. Leakage level: 0.16 bars in 16 hours <= good ;-)Nicolas Chauvin2013-01-11

.

2013-01-11-17:00CETSet helium pressure to 1.00 bar and start calibration with power in nichrome wire.Nicolas Chauvin2013-01-11

.

2013-01-11-21:30CETDetected a bad contact in the power cable for the nichrome wire. And turned off the air flow in the vent hood as it was amplifzing the ocsillation effect of the room temperature control.Nicolas Chauvin2013-01-11

.

2013-01-12-14:45CETPlugged the new T_GlassOutCopper thermocouple into the sensor acquisition of T_Well that we are not using. The temperature measurement with T_GlassOutCopper seems much more stable than our standard T_GlassOut.Nicolas Chauvin2013-01-12

.

2013-01-12-14:50CETT_GlassOut thermocouple replaced in correct position. T_GlassOut thermocouple moved and was not anymore in contact the glass surface. This can explain some the temperature instability with T_GlassOut. But this do not explain the variations with T_GlassIn.Nicolas Chauvin2013-01-12

.

2013-01-12-17:15CETAbort calibration to fix power cable issue.Nicolas Chauvin2013-01-12

.

2013-01-12-18:15CETStart calibration with 1.00 bar of helium and with power in Celani 2L wire.Nicolas Chauvin2013-01-12

.

2013-01-13-13:00CETVacuum the cell and fill it with 1.00 bar of hydrogenNicolas Chauvin2013-01-13

.

2013-01-13-13:40CETStart calibration with 1.00 bar of hydrogen and with power in Celani 2L wire (10W, 20W, 30W)Nicolas Chauvin2013-01-13

.

2013-01-14-10:00CETLoad the cell with 3.1 bar of hydrogen and start loading Celani 2L wire with indirect heatingNicolas Chauvin2013-01-14

.

2013-01-14-14:00CETCelani 2L wire resistivity is not dropping as expected. Try by adding some cycle on the input powerNicolas Chauvin2013-01-14

.

2013-01-14-15:00CETIncrease power to 80 WNicolas Chauvin2013-01-15

.

2013-01-15-08:15CETStop power and let the cell cool downNicolas Chauvin2013-01-15

.

2013-01-15-11:00CETSet hydrogen pressure to 1.00 barNicolas Chauvin2013-01-15

.

2013-01-15-11:20CETStart run with 1.00 bar of hydrogen and 48W in Celani 2L wireNicolas Chauvin2013-01-15

.

2013-01-15-14:00CETIncrease power to 57.2 WNicolas Chauvin2013-01-15

.

2013-01-15-16:00CETIncrease power to 58.6 WNicolas Chauvin2013-01-15

.

2013-01-15-19:30CETStop power and let the cell cool down. Pack the cell to move it for a test in Germany.Nicolas Chauvin2013-01-22

.

2013-01-16-13:30CETInstall cell in test lab in Germany. Hydrogen pressure at 0.992 bar. Set power to 50W.Nicolas Chauvin2013-01-22

.

2013-01-16-18:15CETUnwanted power shutdown of the lab at test facility.Nicolas Chauvin2013-01-22

.

2013-01-17-11:20CETReset measurement system after power down during previous night. Still 1.0 bar of hydrogen. Set power to 50 WNicolas Chauvin2013-01-22

.

2013-01-17-13:30CETPower down due to power surge in the test lab.Nicolas Chauvin2013-01-22

.

2013-01-17-13:45CETReset measurement system and power down again to power surge.Nicolas Chauvin2013-01-22

.

2013-01-17-14:00CETReset measurement system after power surge. Set power to 50 WNicolas Chauvin2013-01-22

.

2013-01-17-15:15CETStop power and let the cell cool down. Pack the cell to move it back to Switzerland lab.Nicolas Chauvin2013-01-22

.

2013-01-21-11:20CETReinstall cell in Swiss lab with still 1.0 bar of hydrogen and set power to 50 W.Nicolas Chauvin2013-01-22

.

2013-01-21-17:20CETIncrease power to 58.6 WNicolas Chauvin2013-01-22

.

2013-01-24-11:30CETHugNetLab laptop disk full. No measurement since 2013-01-23 22:00 CET. Many errors in php scripts filling /var/log/apache2/error.log file. Clean error log files.
PHP script fix needed.
Nicolas Chauvin2013-02-01

.

2013-02-01-14:45CETStop power and let the cell cool down.Nicolas Chauvin2013-02-01

.

2013-02-01-17:30CETRemove Celani 2L wire from the cell. Reassemble the cell with only the nichrome wire in it (blue channel). Set power to 59.0 W on nichrome wire.Nicolas Chauvin2013-02-01

.

2013-02-05-02:15CETProblem with power in Swiss lab. Power for HugNetLab acquisition went down. Power for Internet router also. No data availableNicolas Chauvin2013-02-07

.

2013-02-05-17:00CETFixed lab power problems.Nicolas Chauvin2013-02-07

.

2013-02-07-14:45CETStop power and let the cell cool down.Nicolas Chauvin2013-02-07

.

2013-02-07-15:30CETStop tests. Set Data Interval to Zero. Pack the cell for shipping to FranceNicolas Chauvin2013-02-07

.

2013-02-15-12:00CETCell received from Swizerland.Mathieu Valat2013-02-15

.

2013-02-16-17:00CETMost of the cell is dismantled and prepared for retrofitting to fit with V1.3Mathieu Valat2013-02-15

.

2013-03-08-13:20CETDismentlement has been complete, retrofiting of the original cell is almost done. The new cell V1.3 has been mounted with 2 BLOCK constantan wire in //. Computer is in reconfiguration. The active cell is currently the suject of enviromenent tests, T_amb, R0, Vacuum, TCs...Mathieu Valat2013-03-11

.

2013-03-12-23:10CETCell V1.3A, which the same as the previous one described, is mounted with 2 BLOCK wires, has been loaded with 1bar of He and the power supply has been set to run a calibration the same way to was set for the previous protocol.Mathieu Valat2013-03-12

.

2013-03-22-18:00CETCell V1.3B is prepared for assembly.Mathieu Valat2013-04-03

.

2013-04-02-16:00CETThe dual cell setup has been equiped with a vacuum gauge. Test are going on to verify the level of vacuum it can archieve. The test gave a stabilised value after 1 hour of 2.3mbar with only the first quarter turn valve connected and 2.45mbar with both cells connected.Mathieu Valat2013-04-03

.

2013-04-02-19:00CETCells loaded with 1 bar of He. Calibration script uses setps of 1 hour for 20 to 80W with 5 or 10W steps of P_inMathieu Valat2013-04-03

.

2013-04-03-14:00CETCells under constant vacuum. Calibration script uses steps of 1 hour for 5 to 20W with 5 W stepsMathieu Valat2013-04-03

.

2013-04-05-10:00CETCell A is down because resistance reading spiked when power was in, so regulation in power turned it to maximum under vacuum. Constantan wire sputtered before breaking. Have to make some new mica and change the tube. ETA of back up is Wednesday.Mathieu Valat2013-04-05

.

2013-04-11-16:00CETCell A is back up and reassembled.Mathieu Valat2013-04-12

.

2013-04-12-16:00CETVacuum controller is problematic. It tends to make the reading of current faulty. This has to be rewired. The "one-way" valve is in place and the fastidious tweaking of it is done.Mathieu Valat2013-04-12

.

2013-04-15-17:00CETThe one-way valve is now calibrated correctly and should at the same time hold vacuum in case of electrical power outage and let the vacuum pull 99% of the pump design (2mbar). The dual cell setup is now running calibrations again!Mathieu Valat2013-04-15

.

2013-04-17-14:00CETCarbon sticky tape arrived this morning. It is now mounted on each end of each TC outside the cells. Test are running to see the difference.Mathieu Valat2013-04-17

.

2013-04-24-09:00CETThe regulation inside the room has been improved.Mathieu Valat2013-04-24

.

2013-04-25-13:00CETCalibration using the script of Protocol V2Mathieu Valat2013-05-21

.

2013-05-03-12:00CETLots and lots of improvements on the insulation of the entire Lab. Calibration ETA is Tuesday morning.Mathieu Valat2013-05-03

.

2013-05-13-17:00CETCalibration using the script of Protocol V2Mathieu Valat2013-05-21

.

2013-05-14-09:00CETCalibration using the script of Protocol V2Mathieu Valat2013-05-21

.

2013-05-14-20:20CETCalibration using the script of Protocol V2Mathieu Valat2013-05-21

.

2013-05-15-17:00CETCalibration using the script of Protocol V2Mathieu Valat2013-05-21

.

2013-05-17-12:30CETCalibration using the script of Protocol V2Mathieu Valat2013-05-21

.

2013-05-21-11:00CETCalibration using the script of Protocol V2Mathieu Valat2013-05-21

.

2013-05-23-13:30CETStarted a long run of calibration. 5 times repeated scripts to extract calibration data for phase 0Mathieu Valat2013-05-23

.

2013-06-01-13:45CETCells mounted with identical setup accordingly with Protocol V2. Cell a has A 547L and B 404L coming from the same batch. All tests OK, starting backing of the cells (25W in NiCr for 10ks).Mathieu Valat2013-06-01

.

2013-06-01-13:46CETT_in1 and T_mica are inverted on both cellsMathieu Valat2013-06-01

.

2013-06-05-18:00CETStep 5a, Software issues have been solved. Calibration script is applied to the constantan wiresMathieu Valat2013-06-11

.

2013-06-08-13:30CETStep 5b, Software glitch gave full electrical power in the NiCr wires for less than 2 seconds. NICr are now blue. Calibration script lauched for the monitor wires in both cellsMathieu Valat2013-06-11

.

2013-06-11-10:00CETStep 5c, Blank run: 25W applied on Celani's wires (constantan) to both cells for one weekMathieu Valat2013-06-11

.

2013-06-19-20:39CETToday LOADING of the wires BEGINS! Step 6.1: Cell A is be loaded with 3 bar of H2 and passive heated @30WMathieu Valat2013-06-19

.

2013-06-19-20:40CETThe constantan wire resistance dropped 0.1Ω just after turing the 30W on the heating wire.Mathieu Valat2013-06-19

.

2013-06-19-21:05CETFor a temperature of T_mica=168°C the constantan wire started to drop down its resistanceMathieu Valat2013-06-19

.

2013-06-19-21:07CETAt 169°C the resistance dropped dramaticallyMathieu Valat2013-06-19

.

2013-06-19-21:43CETThe resistance of the wire suddenly increase then dropped down again 5 minutes laterMathieu Valat2013-06-19

.

2013-06-20-09:00CETThe color of the wire has changed. Instead of grayish black it is now coppery looking. T_mica = 169°C INDIRECT HEATING (power to the NiCr)Mathieu Valat2013-06-20

.

2013-06-20-12:00CETTemperatures are rizing. The wire is switching to an other mode. The resistance is still dropping.Mathieu Valat2013-06-20

.

2013-06-20-18:00CETWire was in a different mode of operation. Decided to stop the power in, let it cool down, measure resistance drop and pull vacuum to extract the H2O formed from reduced oxide on the surface of the constantan wire.Mathieu Valat2013-06-20

.

2013-06-20-18:30CETTurning the vacuum pump on triggered the resistance of the constantan wire to drop significantly. This might be due to depressurisation of the gaz leading to cool down the cell.Mathieu Valat2013-06-20

.

2013-06-20-18:40CETResistance drop since the original value is 13.11%Mathieu Valat2013-06-20

.

2013-06-20-19:00CETThe cell A is loaded with 3 bars of H2 and the power input is set to cycle on 7 hours time including 1 hour cooling for 35WMathieu Valat2013-06-20

.

2013-06020-19:30CETThere were power limitation to 31W on the power supply controller. Changed it to 36W and relauchned the setp controllerMathieu Valat2013-06-20

.

2013-06-21-17:40CETDue to personnal occupation I must let the script going on for the rest of the weekend. So the 6 cycles of 35W is relaunched.Mathieu Valat2013-06-21

.

2013-06-23-12:00CETFinal resistance after cooling is 13.11. R/R0 value is 0.757Mathieu Valat2013-06-23

.

2013-06-24:12:00CETCurrents throught Celani wires are optimized for better reading. Cell A is vacuumed. Scripts are ready STEP 7Mathieu Valat2013-06-24

.

2013-06-24-13:30CETBoth contantan wires are wired with their own power supply.Mathieu Valat2013-06-25

.

2013-06-24-21:30CETSTEP 7: Script launched.Mathieu Valat2013-06-25

.

2013-06-25-15:00CETAfter increasing under vacuum heating. The resistance is reaching a plateau above 17ΩMathieu Valat2013-06-25

.

2013-06-25-22:00CETCheck valve remove to have better vacuum speed. Cell A loaded w/ 3b of H2, script will 40W for 36hours indirect heating. PSU set with 40mA contant current for powering Celani wires.Mathieu Valat

.

2013-06-25-22:30CETThe resistance dropped from 17.5 to 14Ω in minutesMathieu Valat

.

2013-06-26-12:20CETPower set to 40W for loading cycle. Current in constantan wires are increased from 40mA to 50mAMathieu Valat

.

2013-06-26-13:40CETSet current of 0.6mA in the Active cell constantan, working on the passive cell to find the right value.Mathieu Valat